OpenStudy (anonymous):

What is the y-intercept of the graph: 3x – 2y = 12 A. (4, 0) B. (4, -6) C. (0, 6) D. (0, -6)

sam (.sam.):

arrange it in slope intercept form 3x – 2y = 12 2y=3x-12 y=3x/2-6

sam (.sam.):

Then y=3x/2-6 y=mx+c c is the y-intercept

sam (.sam.):

We know that when y-intercept , x=0 so, (0, -6)

OpenStudy (anonymous):

\(y\)-intercept when \(x=0\) => \(3(0)-2y=12\) then solve for \(y\)